An executive in an engineering firm earns a monthly salary plus a christmas bonus of $ 8500. If she earn a total of $ 97,3000 per year. What is her monthly salary?
Answer by checkley71(8403)

(97,300-8,500)/12=88,000/12=7,400 for her monthly salary.